If the magnitude of an electric field in air is as great as 3.00×106 N/C, the air becomes ionized and begins to conduct electricity. This phenomenon is called dielectric breakdown. A charge of 14.4 μC is to be placed on a conducting sphere. What is the minimum radius of a sphere that can hold this charge without breakdown?
Answer #1

Electric Field(E) - The region around a charged particle within which a force would be exerted on other charged particles.

Given, E = 3 * 106 N/C

q = 14.4 μC

Lets, consider r to be the radius of the sphere that can hold the charge without breakdown.

Considering a Gaussian Surface in the form of a sphere at radius r , the electric field has the same magnitude at every point of the sphere and is directed outward.The Electric Flux is then just the electric field times the area of the sphere.

Electric Flux = EA = E (4ᴨr2) ........(1)

The total of the electric flux out of a closed surface is equal to the charge enclosed divided by the permittivity

Electric Flux = Q/εo .........................(2)

Combining (1) and (2) we get,

  E (4ᴨr2) =  Q/εo

   r2  =  Q/E(4ᴨεo)

     r2  = 14.4 * 10-6 / (3 * 106 * 4 * 3.14 * 8.85 * 10-12)

= 14.4 / (12 * 3.14 * 8.85)

= 14.4 / 333.468 = 0.04318

r = V0.04318 = 0.2077 m = 20.77 cm

Thus, the minimum radius of the sphere is 20.77 cm